A stationary ball is struck by a bat with an avg force of 60N over 10 ms. If the mass of the ball is .25 kg, what speed does it have just after the impact?
1
Expert's answer
2020-04-09T09:13:32-0400

Solution: The second Newton's law states that the rate of change of momentum of a body is directly proportional to the force applied, and this change in momentum takes place in the direction of the applied force. "\vec F=\frac{d\vec p}{dt}" In our task the directions of force and momentum coincides, and we can write for average magnitudes:

"F=\frac{\Delta p}{\Delta t}" , where increment of the momentum is "\Delta p= m\cdot \Delta v" . Since the ball is initial at rest the speed increment coincides with the final speed "v" of the ball. We have

"F\cdot \Delta t=m \cdot v" , and "v=\frac {F\cdot \Delta t}{m}=\frac{60N\cdot 10\cdot 10^{-3}s}{0.25 kg}=2.4 ms^{-1}"

Answer: speed of the ball just after the impact is 2.4m/s